# Active Preference Learning for Large Language Models

### v1

@@ 5.6 Results
Table 2: Active preference learning results: the mean to 2 d.p. and standard errors to 3 d.p. of the win-rates. For IMDB, we calculate the win rate vs the completions generated by the initial pre-trained model. For TLDR we calculate the win rate vs the human completions available on the test set. The size column represents the size of the acquired dataset used for fine-tuning at the point of evaluation with the forward slash indicating a difference between datasets (IMDB/TLDR)
IMDB TLDR
Size
Random
Pref certainty
Pref + Ent
Random
Pref certainty
Pref + Ent
128
$0.67\pm 0.012$
$\textbf{0.68}\pm 0.003$
$\textbf{0.68}\pm 0.004$
$0.51\pm 0.008$
$0.52\pm 0.012$
$\textbf{0.53}\pm 0.006$
@@
$0.76\pm 0.008$
$\textbf{0.82}\pm 0.005$
$0.81\pm 0.007$
$0.55\pm 0.01$
$\textbf{0.56}\pm 0.012$
$0.55\pm 0.01$
[…]
$0.84\pm 0.004$
$0.87\pm 0.007$
$\textbf{0.89}\pm 0.006$
$0.55\pm 0.009$
$0.58\pm 0.01$
$\textbf{0.59}\pm 0.007$
[…]
$0.88\pm 0.004$
$0.90\pm 0.004$
$\textbf{0.92}\pm 0.005$
$0.56\pm 0.012$
$\textbf{0.60}\pm 0.009$
$0.59\pm 0.01$
@@ 5.6.1 Analysing acquired data
In section 3.1 we motivate why the preference certainty acquisition strategy may provide an advantage versus a random baseline when fine-tuning with DPO. This focused around whether it would surface examples where the implicit preference model provided an incorrect prediction, with certainty. We carry out a post hoc analysis on the data acquired during our experiments to better understand the characteristics of the acquired examples. In particular what differs between the different acquisition strategies and how it changes as fine-tuning phases progress. The approach we take is to look at how the implicit preference predictions from the model correlate with the true oracle preferences.
We construct a classifier using the Bradley Terry (BT) model - equation 6 in (Rafailov et al., 2023) - that gives us $p(y_{1}\succ y_{0}|x)\in[0,1]$ under our implicit reward model (equation 5). Using the probabilities this provides we construct histograms in figure 3 for the batches of $M$ acquired datapoints across all 9 seeds. We map the data in such a way that the bucket at 0.9 will contain examples where the BT model was most confidently correct according to it’s probability, and 0.1 will contain the most confidently wrong. The red $0.0\rightarrow 0.5$ contains all the incorrect predictions bucketed into 10 bins according to their probability. The green $0.5\rightarrow 1$ contains all correct predictions. To determine correctness, we use a 0.5 decision threshold on our BT model and compare the result to the ranking provided by the oracle.
We can see from these histograms that the random acquisition selects quite uniform examples according to the implicit preference model predictions. The preference certainty based acquisition on the other hand surfaces a lot of confidently incorrect examples which will aid fine-tuning with DPO.
Figure 3: Histograms of probabilities from our implicit Bradley Terry preference model across a batch of acquired data; grouped by incorrect (red) and correct (green) preferences according to the oracle. This assumes a decision threshold of 0.5. Our preference certainty acquisition function surfaces confidently wrong examples.
@@ 6 Conclusion and Discussion
We’ve demonstrated a simple and effective way to improve the use of an oracle labelling budget for preference fine-tuning LLMs. Our active learning setup builds upon DPO and uses the implicit preference model to determine which data points to get oracle judgements for during online training.
Given the ever increasing computational cost involved in training SOTA large language models, it is important to consider practical limitations of scaling up our setup. One such example is that we re-initialise the parameters of the model at each fine-tuning step $t$ as done in previous deep active learning works (Gal et al., 2017). This helps us isolate the impact of the different acquisition strategies, which is the focus. A promising direction of future work is to integrate approaches from online learning (Ritter et al., 2018). This could significantly improve computational efficiency by allowing us to not re-initialise the parameters at each time step and spend the majority of the fine-tuning budget on the most recently acquired data. This would involve further changes to the model and/or how we are sampling the data when fine-tuning. An alternative direction here is to explore combining our approach with parameter efficient fine-tuning techniques like LORA (Hu et al., 2021). Acquiring smaller batches with more regular updates would also likely further favour the more active approach.
Another interesting direction of future work is to explore additional data acquisition strategies. For example, we can include measures of diversity of samples within a batch, or taking a more Bayesian approach to explicitly model the epistemic uncertainty of our model (Kirsch et al., 2019).
Lastly, the use of LLMs as evaluators in this setting is of independent interest. Investing more time into getting smaller, more economical models to work for these sorts of use cases would make it easier to run larger amounts of ablations in order to drawer stronger conclusions.
@@
We deal with the problem of fine-tuning large language models. Although the models used in our specific experiments can fit on a single large A100 GPU and are manageable in terms of energy consumption, our framework could be applied to much larger closed source models. This could lead to the indirect negative consequence of this work on the environment, due to the large amount of energy required.
@@ Appendix B Data preprocessing
Prompt
SUBREDDIT: r/cats TITLE: Acquired cat! Now a question.. POST: So, I just got a lovely little cat named Luna. She’s about a year, a year and a half and pretty tiny. I live in an apartment located on the 5th floor of my building. My apartment doesn’t have AC (I’m in NYC) and I usually like to leave the windows open for ventilation. They’ve got child bars, but because Luna is so small she can easily fit through them–and did a few moments ago. Nearly gave me a heart attack watching her slip through them and walked out onto a very narrow ledge 5 floors above a concrete sidewalk. She came right back in, but now I’m concerned about having a dead cat on my hands (or more accurately, on my sidewalk). So my question is, should I trust her cat instincts and leave the windows open? Or shall I sit in a stuffy apartment with the windows sealed? TL;DR:
Human Summary
I live on the 5th floor and my cat just walked out on my window ledge and came back in. Should I be nervous she’s going to explore too far out and fall to her kitty death?
Prompt
SUBREDDIT: r/AskReddit TITLE: I want to throw a great party, Reddit. What are some special things that I can do to make this happen? POST: Alright, so my 19th birthday is next Friday. I live in a college town, renting a house alone for the summer, and I want to throw a great party. I have invited friends from both school (which means they are coming from all over the state) and from home. However, since it’s summer, I know that a lot of people won’t show. Gas is expensive and for some, it’s a far drive, so I understand. I’m thinking music, card playing, beer pong, etc… But what else can I do to make sure that everyone mingles, it isn’t lame, and that things stay under control while everyone still has a good time? TL;DR:
Human Summary
How do I throw a kick ass party with a bunch of people who don’t really know each other and my house is still standing in the morning?
@@ Appendix C Example responses
Prompt
The comments of the previous user are harsh indeed
A: Fine-tuned completion
. This is a great film with a great story and a great cast. Highly recommended. I saw this movie on Sund
B: Original completion
. I am a fan of the original series and I have watched the first two seasons of the series. I have also
Preference
A
Rationale
Review A is better as it expresses a clear positive sentiment towards the film, while Review B does not provide any opinion on the movie
Prompt
I was looking forward to this so much, being a
Fine-tuned completion
huge Harry Potter fan. I was very impressed with the story and the acting. I highly recommend it.
Original completion
huge fan of the original series. I was so disappointed. I was expecting a lot more from this.
Preference
A
Rationale
Review A is better because it provides a positive sentiment about the movie, is grammatically correct, and does not contain repetition like Review B.
@@
Prompt
SUBREDDIT: r/relationships TITLE: I(17M)am gealous because my girlfriend(16F) of over 4 months is having conversation that she starts with my classmate. POST: I am now really down because of this. We had an argument 2 months ago about this and I told her to stay away from my classmate(he is a really bad person and always speaks really bad about her when he is with others or even with me, we’ve been colleagues for 10 years now and he is awful, I hate him), and I’m sick of it. The problem is, today after checking her phone I realized she starts most conversation and he is mostly not answering to her immediately. I am not sure what to do and I am really confused and really really sad and disappointed. What should I do? TL;DR:
A: Fine-tuned completion
GF of over 4 months starts most conversation with my classmate, he is mostly not answering to her immediately. I am not sure what to do and am really sad and disappointed. Need advice. [endoftext]
B: Original completion
I am really sad because my girlfriend starts most conversation with my classmate and I’m sick of it. What should I do?[endoftext]
Preference
A
Rationale
Summary A is preferred as it includes more relevant details from the post, such as the duration of the relationship, the classmate’s lack of immediate response, and the poster’s need for advice, which were missing in Summary B.
Prompt
SUBREDDIT: r/dogs TITLE: Need a little advice regarding our elderly dog. POST: Our rescue [Lt. Ellen Ripley] has been showing her age lately. Actually for the last six months or so she has been going a bit downhill. Formerly very active and happy to play she’s blind, deaf and now sleeps or spins in circles all day. She hasn’t been able to hold her water or poops at all and we’ve been coming home to a nightly cleanup. Worse she’s been biting at us a lot and seems to think our hands are actually delicious fish sticks.
We’ve tried medicine and have seen the vet numerous times but she just seems to be getting worse and worse.
So, advice time. Am I being selfish? I feel like keeping her around may not be the best thing for her anymore. I worry that she spends her time now scared, confused and in the dark. But so far as I know, she’s not in any pain and she can still recognize me on occasion.
I’ll TL;DR:
A: Fine-tuned completion
Old rescue dog has been getting worse and worse. Tried medicine and the vet. Don’t know if keeping her around is the best thing for her anymore. Need advice.[endoftext]
B: Original completion
We have a dog that is getting older and is starting to become more depressed. What should we do about it?[endoftext]
Preference
A
Rationale
Summary A is more accurate and detailed, capturing the severity of the dog’s condition, the attempts at treatment, and the owner’s dilemma, while Summary B oversimplifies the situation by referring to the dog as merely ”more depressed.”
@@
C.1 Fine-tuning iterations

### v2

@@ 5.6 Results
For the first fine-tuning step ($M=128$), there is no discernible difference between the strategies. This makes sense when using the preference certainty acquisition because the initial pre-trained model is used to rank the data and it doesn’t yet know anything about the oracle’s preferences. In Appendix C we provide examples of typical prompt and completion pairs, alongside the oracle preference and rationale provided by our GPT-4 oracle, before and after the fine-tuning process.
Table 2: Active preference learning results: the mean to 2 d.p. and standard errors to 3 d.p. of the win-rates. For IMDB, we calculate the win-rate vs the completions generated by the initial pre-trained model. For TLDR we calculate the win-rate vs the human completions available on the test set. The size column represents the size of the acquired dataset used for fine-tuning at the point of evaluation.
Dataset Size Random Entropy Pref certainty Pref + Ent
IMDB 128 $0.67\pm 0.012$ $0.68\pm 0.011$ $\textbf{0.68}\pm 0.003$ $\textbf{0.68}\pm 0.004$
256 $0.76\pm 0.008$ $0.74\pm 0.009$ $\textbf{0.82}\pm 0.005$ $0.81\pm 0.007$
512 $0.84\pm 0.004$ $0.84\pm 0.009$ $0.87\pm 0.007$ $\textbf{0.89}\pm 0.006$
768 $0.88\pm 0.004$ $0.89\pm 0.004$ $0.90\pm 0.004$ $\textbf{0.92}\pm 0.005$
@@
256 $0.55\pm 0.01$ $\textbf{0.57}\pm 0.01$ $0.56\pm 0.012$ $0.55\pm 0.01$
384 $0.55\pm 0.009$ $0.56\pm 0.017$ $0.58\pm 0.01$ $\textbf{0.59}\pm 0.007$
512 $0.56\pm 0.012$ $0.58\pm 0.01$ $\textbf{0.60}\pm 0.009$ $0.59\pm 0.01$
Figure 3: Histograms of probabilities from our implicit Bradley Terry preference model across a batch of acquired data; grouped by incorrect (red) and correct (green) preferences according to the oracle. This assumes a decision threshold of 0.5. Our preference certainty acquisition function surfaces confidently with wrong examples.
@@ 5.6.1 Analysing acquired data
In section 3.1 we motivate why the preference certainty acquisition strategy may provide an advantage versus a random baseline when fine-tuning with DPO. This focused on whether it would surface examples where the implicit preference model provided an incorrect prediction, with certainty. We carry out a post hoc analysis of the data acquired during our experiments to better understand the characteristics of the acquired examples. In particular, what differs between the different acquisition strategies and how they change as fine-tuning phases progress. The approach we take is to look at how the implicit preference predictions from the model correlate with the true oracle preferences.
We construct a classifier using the Bradley Terry (BT) model - equation 6 in (Rafailov et al., 2023) - that gives us $p(y_{1}\succ y_{0}|x)\in[0,1]$ under our implicit reward model (equation 5). Using the probabilities provided, we construct histograms in figure 3 for the batches of $M$ acquired datapoints across all 9 seeds. We map the data in such a way that the bucket at 0.9 will contain examples where the BT model was most confidently correct according to it’s probability, and 0.1 will contain the most confidently wrong. The red $0.0\rightarrow 0.5$ contains all the incorrect predictions bucketed into 10 bins according to their probability. The green $0.5\rightarrow 1$ contains all correct predictions. To determine correctness, we use a 0.5 decision threshold on our BT model and compare the result to the ranking provided by the oracle.
We can see from these histograms that the random acquisition selects quite uniform examples according to the implicit preference model predictions. The preference certainty-based acquisition on the other hand surfaces a lot of confidently incorrect examples which ultimately aid with improving fine-tuning performance when using DPO.
@@ 6 Conclusion and Discussion
We’ve demonstrated a simple and effective way to improve the use of an oracle labelling budget for preference fine-tuning LLMs. Our active learning setup builds upon DPO and uses the implicit preference model to determine which data points to get oracle judgements during online training.
Given the ever increasing computational cost involved in training SOTA large language models, it is important to consider the practical limitations of scaling up our setup. One such example is that we re-initialise the parameters of the model at each fine-tuning step $t$ as done in previous deep active learning works (Gal et al., 2017). This helps us isolate the impact of the different acquisition strategies, which is the focus. A promising direction of future work is to integrate approaches from online learning (Ritter et al., 2018). This could significantly improve computational efficiency by allowing us to not re-initialise the parameters at each time step and spend the majority of the fine-tuning budget on the most recently acquired data. This could involve further changes to the model and/or how we are sampling the data when fine-tuning. In Appendix D, we discuss minimally adapting our approach for online learning and present promising preliminary results to motivate future work in this direction.
An alternative direction here is to explore combining our approach with parameter-efficient fine-tuning techniques like LORA (Hu et al., 2021). Acquiring smaller batches with more regular updates would also likely further favour the more active approach. Another interesting direction of future work is to explore additional data acquisition strategies. For example, we can include measures of the diversity of samples within a batch, or take a more Bayesian approach to explicitly model the epistemic uncertainty of our model (Kirsch et al., 2019). Lastly, the use of LLMs as evaluators in this setting is of independent interest. Investing more time into getting smaller, more economical models to work for these sorts of use cases would make it easier to run larger amounts of ablations in order to draw stronger conclusions.
@@
We deal with the problem of fine-tuning large language models. Although the models used in our specific experiments can fit on a single large A100 GPU and are manageable in terms of energy consumption, our framework could be applied to much larger closed-source models. This could lead to the indirect negative consequence of this work on the environment, due to the large amount of energy required.
@@ Appendix B Data preprocessing
Prompt SUBREDDIT: r/cats TITLE: Acquired cat! Now a question.. POST: So, I just got a lovely little cat named Luna. She’s about a year, a year and a half and pretty tiny. I live in an apartment located on the 5th floor of my building. My apartment doesn’t have AC (I’m in NYC) and I usually like to leave the windows open for ventilation. They’ve got child bars, but because Luna is so small she can easily fit through them–and did a few moments ago. Nearly gave me a heart attack watching her slip through them and walked out onto a very narrow ledge 5 floors above a concrete sidewalk. She came right back in, but now I’m concerned about having a dead cat on my hands (or more accurately, on my sidewalk). So my question is, should I trust her cat instincts and leave the windows open? Or shall I sit in a stuffy apartment with the windows sealed? TL;DR:
Human Summary I live on the 5th floor and my cat just walked out on my window ledge and came back in. Should I be nervous she’s going to explore too far out and fall to her kitty death?
Prompt SUBREDDIT: r/AskReddit TITLE: I want to throw a great party, Reddit. What are some special things that I can do to make this happen? POST: Alright, so my 19th birthday is next Friday. I live in a college town, renting a house alone for the summer, and I want to throw a great party. I have invited friends from both school (which means they are coming from all over the state) and from home. However, since it’s summer, I know that a lot of people won’t show. Gas is expensive and for some, it’s a far drive, so I understand. I’m thinking music, card playing, beer pong, etc… But what else can I do to make sure that everyone mingles, it isn’t lame, and that things stay under control while everyone still has a good time? TL;DR:
Human Summary How do I throw a kick ass party with a bunch of people who don’t really know each other and my house is still standing in the morning?
@@ Appendix C Example responses
Prompt The comments of the previous user are harsh indeed
A: Fine-tuned completion . This is a great film with a great story and a great cast. Highly recommended. I saw this movie on Sund
B: Original completion . I am a fan of the original series and I have watched the first two seasons of the series. I have also
Preference A
Rationale Review A is better as it expresses a clear positive sentiment towards the film, while Review B does not provide any opinion on the movie
Prompt I was looking forward to this so much, being a
Fine-tuned completion huge Harry Potter fan. I was very impressed with the story and the acting. I highly recommend it.
Original completion huge fan of the original series. I was so disappointed. I was expecting a lot more from this.
Preference A
Rationale Review A is better because it provides a positive sentiment about the movie, is grammatically correct, and does not contain repetition like Review B.
@@
Prompt SUBREDDIT: r/relationships TITLE: I(17M)am gealous because my girlfriend(16F) of over 4 months is having conversation that she starts with my classmate. POST: I am now really down because of this. We had an argument 2 months ago about this and I told her to stay away from my classmate(he is a really bad person and always speaks really bad about her when he is with others or even with me, we’ve been colleagues for 10 years now and he is awful, I hate him), and I’m sick of it. The problem is, today after checking her phone I realized she starts most conversation and he is mostly not answering to her immediately. I am not sure what to do and I am really confused and really really sad and disappointed. What should I do? TL;DR:
A: Fine-tuned completion GF of over 4 months starts most conversation with my classmate, he is mostly not answering to her immediately. I am not sure what to do and am really sad and disappointed. Need advice. [endoftext]
B: Original completion I am really sad because my girlfriend starts most conversation with my classmate and I’m sick of it. What should I do?[endoftext]
Preference A
Rationale Summary A is preferred as it includes more relevant details from the post, such as the duration of the relationship, the classmate’s lack of immediate response, and the poster’s need for advice, which were missing in Summary B.
Prompt SUBREDDIT: r/dogs TITLE: Need a little advice regarding our elderly dog. POST: Our rescue [Lt. Ellen Ripley] has been showing her age lately. Actually for the last six months or so she has been going a bit downhill. Formerly very active and happy to play she’s blind, deaf and now sleeps or spins in circles all day. She hasn’t been able to hold her water or poops at all and we’ve been coming home to a nightly cleanup. Worse she’s been biting at us a lot and seems to think our hands are actually delicious fish sticks. We’ve tried medicine and have seen the vet numerous times but she just seems to be getting worse and worse. So, advice time. Am I being selfish? I feel like keeping her around may not be the best thing for her anymore. I worry that she spends her time now scared, confused and in the dark. But so far as I know, she’s not in any pain and she can still recognize me on occasion. I’ll TL;DR:
A: Fine-tuned completion Old rescue dog has been getting worse and worse. Tried medicine and the vet. Don’t know if keeping her around is the best thing for her anymore. Need advice.[endoftext]
B: Original completion We have a dog that is getting older and is starting to become more depressed. What should we do about it?[endoftext]
Preference A
Rationale Summary A is more accurate and detailed, capturing the severity of the dog’s condition, the attempts at treatment, and the owner’s dilemma, while Summary B oversimplifies the situation by referring to the dog as merely ”more depressed.”
@@
Appendix D Online variation
In section 6 we note that a practical limitation to scaling algorithm 1 to significantly larger models is that the parameters $\theta_{t}$ are reset to $\theta_{0}$ in step 12 and the model is fine-tuned on all the data acquired so far each step. Here we motivate an online variant where we instead don’t reset the parameters and do a single gradient update with respect to the most recently acquired data during fine-tuning. This significantly reduces the time spent fine-tuning during the active learning approach. This is similar to the approach taken in (Guo et al., 2024), where they randomly, instead of actively, acquire the data and consider comparing online vs offline.
We re-run our IMDB experiment from section 5 with this online active variation for random and preference certainty acquisitions. We do 3 random seeds and include standard error bars in figure 5. We find that preference certainty significantly outperforms random on this problem, motivating further study.
Figure 5: Win-rate vs initial model after each acquired batch for IMDB with random and preference certainty acquisition and online fine-tuning. Only a single fine-tuning gradient step is taken on the latest batch.
Appendix E Fine-tuning iterations